# Recursive construction and enumeration of self-orthogonal and self-dual codes over finite commutative chain rings of even characteristic - I

### v1

@@ 2 Some preliminaries
Now, let $\mathcal{C}^{(1)}\subseteq\mathcal{C}^{(2)}\subseteq\cdots\subseteq\mathcal{C}^{(s+\theta_{e})}$ be a chain of self-orthogonal codes of length $n$ over $\mathcal{T}_{m},$ where $\dim\mathcal{C}^{(i)}=\Lambda_{i}$ for $1\leq i\leq s+\theta_{e},$ the code $\mathcal{C}^{(s-\kappa_{1})}$ is doubly even, and the code $\mathcal{C}^{(s-\kappa+\theta_{e})}$ satisfies the additional condition $\textbf{1}\notin\mathcal{C}^{(s-\kappa+\theta_{e})},$ provided that $2\kappa\leq e,$ $n\equiv 4\pmod{8}$ and $m$ is odd. Note that the proofs of Propositions 3.1-3.8 provide a method to construct a self-orthogonal code $\mathscr{D}_{e}$ of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$ satisfying $Tor_{i}(\mathscr{D}_{e})=\mathcal{C}^{(i)}$ for $1\leq i\leq s+\theta_{e}.$ Below, we outline this construction method by distinguishing the following two cases: (A) $2\kappa\leq e$ and (B) $2\kappa>e.$
  1. (A)
    Construction of a self-orthogonal code of length $n$ over $\mathscr{R}_{e,m}$ when $2\kappa\leq e$
    • Input:
      A chain $\mathcal{C}^{(1)}\subseteq\mathcal{C}^{(2)}\subseteq\cdots\subseteq\mathcal{C}^{(s+\theta_{e})}$ of self-orthogonal codes of length $n$ over $\mathcal{T}_{m},$ such that
      • $\bullet$
        $\dim\mathcal{C}^{(i)}=\Lambda_{i}$ for $1\leq i\leq s+\theta_{e},$
      • $\bullet$
        the code $\mathcal{C}^{(s-\kappa_{1})}$ is doubly even, and
      • $\bullet$
        $\textbf{1}\notin\mathcal{C}^{(s-\kappa+\theta_{e})}$ if $n\equiv 4\pmod{8}$ and $m$ is odd.
    • Step 1:
      If $e$ is even, use Proposition 3.1 to construct a self-orthogonal code $\mathscr{D}_{2}$ of type $\{\Lambda_{s},\lambda_{s+1}\}$ and length $n$ over $\mathscr{R}_{2,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{2})=\mathcal{C}^{(s)}.$
      On the other hand, if $e$ is odd, employ Proposition 3.2 to construct a self-orthogonal code $\mathscr{D}_{3}$ of type $\{\Lambda_{s},\lambda_{s+1},\lambda_{s+2}\}$ and length $n$ over $\mathscr{R}_{3,m}$ satisfying property $(\mathfrak{P}),$ $Tor_{1}(\mathscr{D}_{3})=\mathcal{C}^{(s)}$ and $Tor_{2}(\mathscr{D}_{3})=\mathcal{C}^{(s+1)}.$
    • Step 2:
      For $4\leq\ell\leq\kappa$ and $\ell\equiv e\pmod{2}$, use Proposition 3.3 to construct a self-orthogonal code $\mathscr{D}_{\ell}$ of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and length $n$ over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\ell})=\mathcal{C}^{(\gamma_{\ell}+1)}$ and $Tor_{j+1}(\mathscr{D}_{\ell})=Tor_{j}(\mathscr{D}_{\ell-2})$ for $1\leq j\leq\ell-2$ from a self-orthogonal code $\mathscr{D}_{\ell-2}$ of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P}).$
    • Step 3:
      When $e$ is even, employ Proposition 3.4 to construct a self-orthogonal code $\mathscr{D}_{\kappa+1}$ of type $\{\Lambda_{s-\kappa_{1}},\\ \lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+1}\}$ and length $n$ over $\mathscr{R}_{\kappa+1,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\kappa+1})=\mathcal{C}^{(s-\kappa_{1})}$ and $Tor_{j+1}(\mathscr{D}_{\kappa+1})=Tor_{j}(\mathscr{D}_{\kappa-1})$ for $1\leq j\leq\kappa-1$ from a self-orthogonal code $\mathscr{D}_{\kappa-1}$ of type $\{\Lambda_{s-\kappa_{1}+1},\lambda_{s-\kappa_{1}+2},\ldots,\lambda_{s+\kappa_{1}}\}$ and length $n$ over $\mathscr{R}_{\kappa-1,m}$ satisfying property $(\mathfrak{P}).$
      On the other hand, when $e$ is odd, use Proposition 3.5 to construct a self-orthogonal code $\mathscr{D}_{\kappa+2}$ of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+2}\}$ and length $n$ over $\mathscr{R}_{\kappa+2,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\kappa+2})=\mathcal{C}^{(s-\kappa_{1})}$ and $Tor_{j+1}(\mathscr{D}_{\kappa+2})=Tor_{j}(\mathscr{D}_{\kappa})$ for $1\leq j\leq\kappa$ from a self-orthogonal code $\mathscr{D}_{\kappa}$ of type $\{\Lambda_{s-\kappa_{1}+1},\lambda_{s-\kappa_{1}+2},\ldots,\lambda_{s+\kappa_{1}+1}\}$ and length $n$ over $\mathscr{R}_{\kappa,m}$ satisfying property $(\mathfrak{P}).$
    • Step 4:
      For $\kappa+3\leq\ell\leq e-\kappa+1$ and $\ell\equiv e\pmod{2}$, use Proposition 3.6 to construct a self-orthogonal code $\mathscr{D}_{\ell}$ of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and length $n$ over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\ell})=\mathcal{C}^{(\gamma_{\ell}+1)}$ and $Tor_{i+1}(\mathscr{D}_{\ell})=Tor_{i}(\mathscr{D}_{\ell-2})$ for $1\leq i\leq\ell-2$ from a self-orthogonal code $\mathscr{D}_{\ell-2}$ of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P}).$
    • Step 5:
      For $e-\kappa+2\leq\ell\leq e$ and $\ell\equiv e\pmod{2}$, employ Proposition 3.8 to construct a self-orthogonal code $\mathscr{D}_{\ell}$ of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and length $n$ over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\ell})=\mathcal{C}^{(\gamma_{\ell}+1)}$ and $Tor_{i+1}(\mathscr{D}_{\ell})=Tor_{i}(\mathscr{D}_{\ell-2})$ for $1\leq i\leq\ell-2$ from a self-orthogonal code $\mathscr{D}_{\ell-2}$ of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P}).$
    • Output:
      A self-orthogonal code $\mathscr{D}_{e}$ of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$ satisfying $Tor_{i}(\mathscr{D}_{e})=\mathcal{C}^{(i)}$ for $1\leq i\leq s+\theta_{e}.$
  2. (B)
    Construction of a self-orthogonal code of length $n$ over $\mathscr{R}_{e,m}$ when $2\kappa>e.$
    • Input:
      A chain $\mathcal{C}^{(1)}\subseteq\mathcal{C}^{(2)}\subseteq\cdots\subseteq\mathcal{C}^{(s+\theta_{e})}$ of self-orthogonal codes of length $n$ over $\mathcal{T}_{m},$ where $\dim\mathcal{C}^{(i)}=\Lambda_{i}$ for $1\leq i\leq s+\theta_{e},$ and the code $\mathcal{C}^{(s-\kappa_{1})}$ is doubly even.
    • Step 1:
      If $e$ is even, apply Proposition 3.1 to construct a self-orthogonal code $\mathscr{D}_{2}$ of type $\{\Lambda_{s},\lambda_{s+1}\}$ and length $n$ over $\mathscr{R}_{2,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{2})=\mathcal{C}^{(s)}.$
      On the other hand, if $e$ is odd, apply Proposition 3.2 to construct a self-orthogonal code $\mathscr{D}_{3}$ of type $\{\Lambda_{s},\lambda_{s+1},\lambda_{s+2}\}$ and length $n$ over $\mathscr{R}_{3,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{3})=\mathcal{C}^{(s)}$ and $Tor_{2}(\mathscr{D}_{3})=\mathcal{C}^{(s+1)}.$
    • Step 2:
      For $4\leq\ell\leq e-\kappa+1-2\theta_{e}$ and $\ell\equiv e\pmod{2}$, apply Proposition 3.3 to construct a self-orthogonal code $\mathscr{D}_{\ell}$ of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and length $n$ over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\ell})=\mathcal{C}^{(\gamma_{\ell}+1)}$ and $Tor_{j+1}(\mathscr{D}_{\ell})=Tor_{j}(\mathscr{D}_{\ell-2})$ for $1\leq j\leq\ell-2$ from a self-orthogonal code $\mathscr{D}_{\ell-2}$ of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P}).$
    • Step 3:
      For $e-\kappa+2-2\theta_{e}\leq\ell\leq\kappa-\mathrm{f}_{2\kappa-e}+1$ and $\ell\equiv e\pmod{2}$, apply Proposition 3.7 to construct a self-orthogonal code $\mathscr{D}_{\ell}$ of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and length $n$ over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\ell})=\mathcal{C}^{(\gamma_{\ell}+1)}$ and $Tor_{i+1}(\mathscr{D}_{\ell})=Tor_{i}(\mathscr{D}_{\ell-2})$ for $1\leq i\leq\ell-2$ from a self-orthogonal code $\mathscr{D}_{\ell-2}$ of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P}).$
    • Step 4:
      For $\kappa-\mathrm{f}_{2\kappa-e}+1<\ell\leq e$ and $\ell\equiv e\pmod{2}$, apply Proposition 3.8 to construct a self-orthogonal code $\mathscr{D}_{\ell}$ of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and length $n$ over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\ell})=\mathcal{C}^{(\gamma_{\ell}+1)}$ and $Tor_{i+1}(\mathscr{D}_{\ell})=Tor_{i}(\mathscr{D}_{\ell-2})$ for $1\leq i\leq\ell-2$ from a self-orthogonal code $\mathscr{D}_{\ell-2}$ of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P}).$
    • Output:
      A self-orthogonal code $\mathscr{D}_{e}$ of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$ satisfying $Tor_{i}(\mathscr{D}_{e})=\mathcal{C}^{(i)}$ for $1\leq i\leq s+\theta_{e}.$
@@
When $\lambda_{j}=\lambda_{e-j+2}$ for $1\leq j\leq e,$ we see, by Lemma 2.1, that the above methods give rise to construction of a self-dual code $\mathscr{D}_{e}$ of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$ satisfying $Tor_{i}(\mathscr{D}_{e})=\mathcal{C}^{(i)}$ for $1\leq i\leq s+\theta_{e}.$ Additionally, the recursive construction method described above coincides with the one used in [32] when $\mathfrak{s}=1.$ In the following example, we illustrate that the self-orthogonal code $\mathcal{C}_{0},$ as defined in Example 2.1, can be lifted to a self-orthogonal code over $\mathscr{R}_{8,2}$ by applying the above recursive construction method (A), as $2\kappa=6<8=e$ in this case.
@@
Here, we will show that the code $\mathcal{C}_{0}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ and length $4$ over $\mathscr{R}_{8,2}$ by applying the above recursive construction method (A). Towards this, let us consider a linear code $\mathscr{C}_{2}$ of type $\{1,0\}$ and length $4$ over $\mathscr{R}_{2,2}$ with a generator matrix
$\mathcal{G}_{2}=\begin{bmatrix}1&1&1&1\end{bmatrix}+u\begin{bmatrix}0&a_{1}&b_{1}&c_{1}\end{bmatrix}$,
where $a_{1},b_{1},c_{1}\in\mathcal{T}_{2}.$ Now, by the recursive construction method (A), we see that the code $\mathcal{C}_{0}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ over $\mathscr{R}_{8,2}$ if and only if $\mathscr{C}_{2}$ is a self-orthogonal code over $\mathscr{R}_{2,2}$ satisfying property $(\mathfrak{P}),$ which holds if and only if $a_{1}^{2}+b_{1}^{2}+c_{1}^{2}\equiv 0\pmod{u^{2}}$, or equivalently, $a_{1}+b_{1}+c_{1}\equiv 0\pmod{u}$ holds. It is easy to see that $a_{1},b_{1},c_{1}\in\mathcal{T}_{2}$ satisfying $a_{1}+b_{1}+c_{1}\equiv 0\pmod{u}$ have precisely $16$ distinct choices. Thus, for a fixed choice of $a_{1},b_{1},c_{1}\in\mathcal{T}_{2}$ satisfying $a_{1}+b_{1}+c_{1}\equiv 0\pmod{u}$, the code $\mathscr{C}_{2}$ of length $4$ over $\mathscr{R}_{2,2}$ with a generator matrix $\mathcal{G}_{2}$ is a self-orthogonal code satisfying property $(\mathfrak{P}).$ Further, note that $\mathcal{G}_{2}\mathcal{G}_{2}^{t}\equiv u^{2}(a_{1}^{2}+b_{1}^{2}+c_{1}^{2})+u^{4}(a_{1}+b_{1}+c_{1})+u^{6}\pmod{u^{7}}$. Since $a_{1}+b_{1}+c_{1}\equiv 0\pmod{u}$, let us suppose that $a_{1}+b_{1}+c_{1}\equiv u\gamma_{1}+u^{2}\gamma_{2}+u^{3}\gamma_{3}+u^{4}\gamma_{4}+u^{5}\gamma_{5}+u^{6}\gamma_{6}\pmod{u^{7}}$, where $\gamma_{i}\in\mathcal{T}_{2}$ for $1\leq i\leq 6.$ This implies that $\mathcal{G}_{2}\mathcal{G}_{2}^{t}\equiv u^{4}\gamma_{1}^{2}+u^{5}\gamma_{1}+u^{6}(1+\gamma_{2}+\gamma_{2}^{2})\pmod{u^{7}}$.
@@
$\mathcal{G}_{4}=\mathcal{G}_{2}+u^{2}\begin{bmatrix}0&a_{2}&b_{2}&c_{2}\end{bmatrix}+u^{3}\begin{bmatrix}0&a_{3}&b_{3}&c_{3}\end{bmatrix}$,
where $a_{2},b_{2},c_{2},a_{3},b_{3},c_{3}\in\mathcal{T}_{2}.$ Using again the recursive construction method (A), we see that the code $\mathscr{C}_{4}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ over $\mathscr{R}_{8,2}$ if and only if $\mathscr{C}_{4}$ is a self-orthogonal code over $\mathscr{R}_{4,2}$ satisfying property $(\mathfrak{P}),$ which holds if and only if $\mathcal{G}_{4}\mathcal{G}_{4}^{t}\equiv 0\pmod{u^{7}}$. This holds if and only if $\gamma_{1}^{2}+a_{2}^{2}+b_{2}^{2}+c_{2}^{2}+u(\gamma_{1}+a_{2}+b_{2}+c_{2})+u^{2}(1+\gamma_{2}+\gamma_{2}^{2}+a_{3}^{2}+b_{3}^{2}+c_{3}^{2}+a_{3}+b_{3}+c_{3})\equiv 0\pmod{u^{3}}$. Let us first choose $a_{2},b_{2},c_{2}\in\mathcal{T}_{2}$ satisfying $a_{2}+b_{2}+c_{2}+\gamma_{1}\equiv 0\pmod{u}$, (note that $a_{2},b_{2},c_{2}$ have precisely $16$ distinct choices). Next, for a fixed choice of $a_{2},b_{2},c_{2}\in\mathcal{T}_{2}$ satisfying $a_{2}+b_{2}+c_{2}+\gamma_{1}\equiv 0\pmod{u}$, let us suppose that $a_{2}+b_{2}+c_{2}+\gamma_{1}\equiv uz_{1}+u^{2}z_{2}\pmod{u^{3}}$. Thus, we need to choose $a_{3},b_{3},c_{3}$ satisfying
$1+\gamma_{2}+\gamma_{2}^{2}+a_{3}^{2}+b_{3}^{2}+c_{3}^{2}+a_{3}+b_{3}+c_{3}+z_{1}+z_{1}^{2}\equiv 0\pmod{u}$. (3.50)
@@
$\mathcal{G}_{6}=\mathcal{G}_{4}+u^{4}\begin{bmatrix}0&a_{4}&b_{4}&c_{4}\end{bmatrix}+u^{5}\begin{bmatrix}0&a_{5}&b_{5}&c_{5}\end{bmatrix}$,
where $a_{4},b_{4},c_{4},a_{5},b_{5},c_{5}\in\mathcal{T}_{2}.$ By the recursive construction method (A), we see that the code $\mathscr{C}_{6}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ over $\mathscr{R}_{8,2}$ if and only if $\mathscr{C}_{6}$ is a self-orthogonal code over $\mathscr{R}_{6,2}$ satisfying property $(\mathfrak{P}),$ which holds if and only if $w+a_{4}+b_{4}+c_{4}\equiv 0\pmod{u}$. It is easy to see that $a_{4},b_{4},c_{4},a_{5},b_{5},c_{5}\in\mathcal{T}_{2}$ satisfying $w+a_{4}+b_{4}+c_{4}\equiv 0\pmod{u}$ have precisely $1024$ distinct choices. Thus, for a fixed choice of $a_{4},b_{4},c_{4},a_{5},b_{5},c_{5}\in\mathcal{T}_{2}$ satisfying $w+a_{4}+b_{4}+c_{4}\equiv 0\pmod{u}$, we see that the code $\mathscr{C}_{6}$ of length $4$ over $\mathscr{R}_{6,2}$ with a generator matrix $\mathcal{G}_{6}$ is a self-orthogonal code over $\mathscr{R}_{6,2}$ satisfying property $(\mathfrak{P}).$
@@
where $a_{6},b_{6},c_{6},a_{7},b_{7},c_{7}\in\mathcal{T}_{2}.$ Note that the code $\mathscr{C}_{8}$ is self-orthogonal for all choices of $a_{6},b_{6},c_{6},a_{7},b_{7},c_{7}\in\mathcal{T}_{2}.$ This shows that the code $\mathcal{C}_{0}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ and length $4$ over $\mathscr{R}_{8,2}$ with the help of the recursive construction method (A).

### v2

@@ 2.6 Notations and preliminary setup
Now, let $\mathcal{C}^{(1)}\subseteq\mathcal{C}^{(2)}\subseteq\cdots\subseteq\mathcal{C}^{(s+\theta_{e})}$ be a chain of self-orthogonal codes of length $n$ over $\mathcal{T}_{m},$ where $\dim\mathcal{C}^{(i)}=\Lambda_{i}$ for $1\leq i\leq s+\theta_{e},$ the code $\mathcal{C}^{(s-\kappa_{1})}$ is doubly even, and the code $\mathcal{C}^{(s-\kappa+\theta_{e})}$ satisfies the additional condition $\textbf{1}\notin\mathcal{C}^{(s-\kappa+\theta_{e})},$ provided that $2\kappa\leq e,$ $n\equiv 4\pmod{8}$ and $m$ is odd. Note that the proofs of Propositions 3.13.5 provide a method to construct a self-orthogonal code $\mathscr{D}_{e}$ of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$ satisfying $Tor_{i}(\mathscr{D}_{e})=\mathcal{C}^{(i)}$ for $1\leq i\leq s+\theta_{e}.$ Figures 1 and 2 present flowcharts outlining the construction method, distinguishing between the two following cases: (I) $2\kappa\leq e,$ and (II) $2\kappa>e.$
@@
When $\lambda_{j}=\lambda_{e-j+2}$ for $1\leq j\leq e,$ we see, by Lemma 2.1, that the construction methods outlined in Figures 1 and 2 give rise to a self-dual code $\mathscr{D}_{e}$ of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m},$ satisfying $Tor_{i}(\mathscr{D}_{e})=\mathcal{C}^{(i)}$ for $1\leq i\leq s+\theta_{e},$ in the cases $2\kappa\leq e$ and $2\kappa>e,$ respectively. Additionally, the recursive construction methods coincide with that used in [35] when $\mathfrak{s}=1.$
In the following example, we illustrate that the self-orthogonal code $\mathcal{C}_{0},$ as defined in Example 2.1, can be lifted to a self-orthogonal code over $\mathscr{R}_{8,2}$ by applying the recursive construction method outlined in Figure 1, since $2\kappa=6<8=e$ in this case.
@@
Here, we will show that the code $\mathcal{C}_{0}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ and length $4$ over $\mathscr{R}_{8,2}$ by applying the recursive construction method outlined in Figure 1. Towards this, let us consider a linear code $\mathscr{C}_{2}$ of type $\{1,0\}$ and length $4$ over $\mathscr{R}_{2,2}$ with a generator matrix
$\vskip-4.2679pt\mathcal{G}_{2}=\begin{bmatrix}1&1&1&1\end{bmatrix}+u\begin{bmatrix}0&a_{1}&b_{1}&c_{1}\end{bmatrix}$,
where $a_{1},b_{1},c_{1}\in\mathcal{T}_{2}.$ Now, by the recursive construction method outlined in Figure 1, we see that the code $\mathcal{C}_{0}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ over $\mathscr{R}_{8,2}$ if and only if $\mathscr{C}_{2}$ is a self-orthogonal code over $\mathscr{R}_{2,2}$ satisfying property $(\mathfrak{P}),$ which holds if and only if $a_{1}^{2}+b_{1}^{2}+c_{1}^{2}\equiv 0\pmod{u^{2}}$, or equivalently, $a_{1}+b_{1}+c_{1}\equiv 0\pmod{u}$ holds. It is easy to see that $a_{1},b_{1},c_{1}\in\mathcal{T}_{2}$ satisfying $a_{1}+b_{1}+c_{1}\equiv 0\pmod{u}$ have precisely $16$ distinct choices. Thus, for a fixed choice of $a_{1},b_{1},c_{1}\in\mathcal{T}_{2}$ satisfying $a_{1}+b_{1}+c_{1}\equiv 0\pmod{u}$, the code $\mathscr{C}_{2}$ of length $4$ over $\mathscr{R}_{2,2}$ with a generator matrix $\mathcal{G}_{2}$ is a self-orthogonal code satisfying property $(\mathfrak{P}).$ Further, note that $\mathcal{G}_{2}\mathcal{G}_{2}^{t}\equiv u^{2}(a_{1}^{2}+b_{1}^{2}+c_{1}^{2})+u^{4}(a_{1}+b_{1}+c_{1})+u^{6}\pmod{u^{7}}$. Since $a_{1}+b_{1}+c_{1}\equiv 0\pmod{u}$, let us suppose that $a_{1}+b_{1}+c_{1}\equiv u\gamma_{1}+u^{2}\gamma_{2}+u^{3}\gamma_{3}+u^{4}\gamma_{4}+u^{5}\gamma_{5}+u^{6}\gamma_{6}\pmod{u^{7}}$, where $\gamma_{i}\in\mathcal{T}_{2}$ for $1\leq i\leq 6.$ This implies that $\mathcal{G}_{2}\mathcal{G}_{2}^{t}\equiv u^{4}\gamma_{1}^{2}+u^{5}\gamma_{1}+u^{6}(1+\gamma_{2}+\gamma_{2}^{2})\pmod{u^{7}}$.
@@
$\vskip-4.2679pt\mathcal{G}_{4}=\mathcal{G}_{2}+u^{2}\begin{bmatrix}0&a_{2}&b_{2}&c_{2}\end{bmatrix}+u^{3}\begin{bmatrix}0&a_{3}&b_{3}&c_{3}\end{bmatrix}$,
where $a_{2},b_{2},c_{2},a_{3},b_{3},c_{3}\in\mathcal{T}_{2}.$ Using again the recursive construction method outlined in Figure 1, we see that the code $\mathscr{C}_{4}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ over $\mathscr{R}_{8,2}$ if and only if $\mathscr{C}_{4}$ is a self-orthogonal code over $\mathscr{R}_{4,2}$ satisfying property $(\mathfrak{P}),$ which holds if and only if $\mathcal{G}_{4}\mathcal{G}_{4}^{t}\equiv 0\pmod{u^{7}}$. This holds if and only if $\gamma_{1}^{2}+a_{2}^{2}+b_{2}^{2}+c_{2}^{2}+u(\gamma_{1}+a_{2}+b_{2}+c_{2})+u^{2}(1+\gamma_{2}+\gamma_{2}^{2}+a_{3}^{2}+b_{3}^{2}+c_{3}^{2}+a_{3}+b_{3}+c_{3})\equiv 0\pmod{u^{3}}$. Let us first choose $a_{2},b_{2},c_{2}\in\mathcal{T}_{2}$ satisfying $a_{2}+b_{2}+c_{2}+\gamma_{1}\equiv 0\pmod{u}$, (note that $a_{2},b_{2},c_{2}$ have precisely $16$ distinct choices). Next, for a fixed choice of $a_{2},b_{2},c_{2}\in\mathcal{T}_{2}$ satisfying $a_{2}+b_{2}+c_{2}+\gamma_{1}\equiv 0\pmod{u}$, let us suppose that $a_{2}+b_{2}+c_{2}+\gamma_{1}\equiv uz_{1}+u^{2}z_{2}\pmod{u^{3}}$. Thus, we need to choose $a_{3},b_{3},c_{3}$ satisfying
[…]
@@
[…]
where $a_{4},b_{4},c_{4},a_{5},b_{5},c_{5}\in\mathcal{T}_{2}.$ By the recursive construction method outlined in Figure 1, we see that the code $\mathscr{C}_{6}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ over $\mathscr{R}_{8,2}$ if and only if $\mathscr{C}_{6}$ is a self-orthogonal code over $\mathscr{R}_{6,2}$ satisfying property $(\mathfrak{P}),$ which holds if and only if $w+a_{4}+b_{4}+c_{4}\equiv 0\pmod{u}$. It is easy to see that $a_{4},b_{4},c_{4},a_{5},b_{5},c_{5}\in\mathcal{T}_{2}$ satisfying $w+a_{4}+b_{4}+c_{4}\equiv 0\pmod{u}$ have precisely $1024$ distinct choices. Thus, for a fixed choice of $a_{4},b_{4},c_{4},a_{5},b_{5},c_{5}\in\mathcal{T}_{2}$ satisfying $w+a_{4}+b_{4}+c_{4}\equiv 0\pmod{u}$, we see that the code $\mathscr{C}_{6}$ of length $4$ over $\mathscr{R}_{6,2}$ with a generator matrix $\mathcal{G}_{6}$ is a self-orthogonal code over $\mathscr{R}_{6,2}$ satisfying property $(\mathfrak{P}).$
@@
where $a_{6},b_{6},c_{6},a_{7},b_{7},c_{7}\in\mathcal{T}_{2}.$ Note that the code $\mathscr{C}_{8}$ is self-orthogonal for all choices of $a_{6},b_{6},c_{6},a_{7},b_{7},c_{7}\in\mathcal{T}_{2}.$ This shows that the code $\mathcal{C}_{0}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ and length $4$ over $\mathscr{R}_{8,2}$ with the help of the recursive construction method described in Figure 1.